### v1

# Towards scalable robotic intervention of children with Autism Spectrum Disorder using LLMs

Abstract
In this paper, we propose a social robot capable of verbally interacting with children with Autism Spectrum Disorder (ASD). This communication is meant to teach perspective-taking using text generated using a Large Language Model (LLM) pipeline. The social robot NAO acts as a stimulator (verbally describes a social situation and asks a question), prompter (presents three options to choose from), and reinforcer (praises when the answer is correct). For the role of the stimulator, the social situation, questions, and options are generated using our LLM pipeline. We compare two approaches: GPT-2 + BART and GPT-2 + GPT-2, where the first GPT-2 common between the pipelines is used for unsupervised social situation generation. We use the SOCIALIQA dataset to fine-tune all of our LLM pipelines. We found that the GPT-2 + BART pipeline had a better BERTscore for generating the questions and the options by combining their individual loss functions. This observation was also consistent with the human evaluations. Lastly, the unsupervised generation of social situations was visualized using T-SNE plots, and the entire pipeline was evaluated for appropriateness for children with ASD by human experts.
Index Terms:
Social HRI, AI-Enabled Robotics, Autonomous AgentsI Introduction
Social robots have had an immense impact on the lives of people in the past decade [1]. They currently share various scenarios of coexistence with humans. This coexistence accounts for Human-Robot Interaction (HRI) in domestic environments like the example of the Roomba robot [2] or of a more interactive robot like Autom that was built to make behavior changes in a more naturalistic user setting [3].
Within healthcare, social robots have undertaken diverse roles too. These roles have required the robot to be involved in a multi-modal interaction with humans. These include engaging in multi-modal interactions with humans, as exemplified by the telepresence robot discussed in [4], facilitating both visual and verbal communication. Other examples of robots being used in healthcare include robots used in surgery [5, 6]. Moreover, robots have found application in psychological interventions, representing another domain where their presence can be immensely notable, as has been seen in ASD interventions [7]. [8].
Autism Spectrum Disorder, as described by the DSM-5, describes individuals that exhibit: 1) challenges in social communication and interaction and 2) restricted and repetitive behaviors [9]. Since there is no cure for ASD, there are countless interventions and treatments, as has been highlighted by the authors in [10]. As robotic intervention of ASD is getting popular with evidences of positive outcomes [11], the development and evaluation of such platforms have gained momentum. Since the population of people with ASD is 1.5%, with every 1 of 54 children in the US alone, sustainability in robotic intervention becomes essential.
Since children with ASD lack an understanding of social cues and situations, a robot that can teach such skills has been a widely proposed solution [12]. So far, the solutions in the literature for robotic intervention have had limited scope for scalability because of the number of finite responses, actions, or pre-defined contents [13] that the robot can generate. Hence, there is a need to bridge this gap with the natural language processing literature that can be used to generate text and speech which the robot can deliver specifically for children with ASD to help them understand various social skills.
To the best of our knowledge, this is the first work in the literature that attempts to increase the scalability of robotic intervention solutions by generating the contents autonomously to teach perspective-taking to children with ASD in a therapy session. In this paper, we make an effort to bridge this gap between language models and their embodiment as a robot for an intervention of children with ASD. In addition, the robot also takes up the role of a prompter and reinforcer (described in Section IV).
This paper has been arranged in the following way: Section II outlines the existing work in the literature related to robots in ASD intervention (Section II-A), robots to teach perspective-taking (II-B), and text generation (Section II-C). Section III highlights the primary contribution of the paper followed by Section IV which describes the workflow highlighting how the robot acts as a stimulator, prompter and reinforcer. Further, Section V describes the problem formulation that we use in this paper. Section VI outlines the results and discussions culminating in Section VII describing the conclusion.
II Related Work
II-A Robots in ASD intervention
The robotic intervention of children with Autism Spectrum Disorder has witnessed variety in both the choice of robot and the respective therapeutic or educational setting. This approach has facilitated the integration of classical methodologies combined with a socially assistive robot, which has urged researchers to examine its effectiveness. In this direction, the authors in [14] have evaluated a Wizard-of-Oz style therapy following the Applied Behavior Analysis (ABA) structure. Based on the efficacy metrics proposed (skill execution and prompt dependency), they could make better conclusions on the clinical merits of a robot-assisted intervention of ASD, which contradicts the HRI metrics that considers gaze, communication, and affect. They also demonstrated the positive effects of such an intervention on one of their study participants. Similar studies done by the authors in [15] have also contributed to positive conclusions about social robots.
The authors in [16] have tested their hypothesis of positive outcomes in a Pivotal Response Treatment (PRT) with robot assistance. They reported that robotic assistance in PRT had a positive effect at the end of the session and showed likeability towards the robot. Furthermore, the results obtained in [17] pointed towards the generalization of improvement in the sociability of the participants in non-robot scenarios. In [18], a new low-cost platform was field tested to see the efficacy of a robot-assisted intervention to improve spontaneous speech, communication, and social skills. Based on statistical analysis, the authors could compare between a robot- and no-robot-assisted groups and could claim relatively higher growth in the former scenario. In addition to research articles, several review articles have pointed towards the productivity of using robot-assisted interventions for individuals with ASD [19, 20, 21].
However, there are limitations to proving an absolute advantage of using robot-assisted intervention compared to a human-lead session [22]. Although both of these sessions have proven beneficial to fetch positive outcomes, robots, compared to humans, have yet to be considered serious agents owing to the lack of sophistication they can offer [23, 22]. Some of these limitations can be attributed to the type of robot employed in a therapy scenario [23]. Since different kinds of robots have been used so far in various intervention scenarios [21, 24, 25], there is no robot-specific intervention protocol mentioned in the Evidence Based Practices for the ‘Technology-Aided Instruction and Intervention (TAII)’ [11]. Although in [26], a more robot-specific scenario has been described by the authors using the KASPER robot for underlining the roles it can take up in a therapy session, it still possesses limitations in the finite reaction possibilities the robot can generate. The robot’s finite response generation ability also adds to the effort a therapist has to put in during a therapy session. As pointed out in [27], the therapist’s involvement in generating content or controlling the robot, as in a Wizard-of-Oz approach, is not sustainable in the long run. Hence, a more sustainable autonomous solution is needed in which the robot can take up the role of a stimulator that can generate and deliver content during a therapy session [26, 27].
II-B Robots to teach perspective-taking and theory of mind
Theory of mind (ToM), as defined in [28], is the ability to understand another person’s mental state to aid the understanding of their social scenario (perspective-taking). Unlike neurotypical children, children with ASD have difficulties perceiving how another person might react in a particular social situation or what personal thoughts might arise out of them in such scenarios [29, 30, 31]. In order to enhance the understanding of these social scenarios, many approaches have been adopted in the ToM interventions.
The authors in [32] have employed a combination of tasks, including learning different scenarios through visual and verbal modalities, to understand perspective-taking. Similarly, in [33], verbal communication and graphical cues-based approach have been followed to assess the beliefs of an individual evaluated based on social scenarios.
Human therapists were involved throughout the session for both of the above ToM-based interventions. Unlike these approaches, the authors in [34, 13] have used a social robot to teach social skills aligned towards the ToM and perspective-taking realm. However, in these studies, the robots followed a pre-defined script of verbal communication to deliver the contents.
II-C Content generation using Language Models for ASD intervention
Text generation, as such, is not a new problem in the natural language processing (NLP) literature. Recently, language models have seen a boom after the introduction of the transformers [35] architecture on various NLP tasks. They have outperformed their RNN-based counterparts for machine translation, text generation, etc. [36, 37] because RNNs suffer from the vanishing gradient problems because of their sequential nature [38, 39]. On the other hand, transformers use a self-attention mechanism, which provides a more parallel computational scheme that is more effective in generating long textual sequences [40].
There can be three broad categories to classify Large Language Models (LLMs) (based on the transformers architecture) for text generation. These include the encoder-only architectures like Bidirectional Encoder Representations from Transformers (BERT) [41] where beam search can be used to collect the most probable subsequent tokens to predict in a sequential manner [42, 43]. The second type of model is the decoder-only architectures like OpenAI’s Generative Pre-trained Transformer (GPT), which has immensely been used for text generation across different fields of study [44, 45, 46]. The third category comprises a model with an encoder-decoder architecture like the BART model, which is used for tasks like text summarization, classification, etc. [47].
Recently, LLMs have been used for robotic specific tasks [48, 49, 50, 51]. This also involves HRI scenarios where goal-oriented LLM usage has been embodied as a robot. Examples include robots being used as a teaching assistant [52], involved in task planning [53] or, being used as an intervention tool in healthcare [54].
Further, generating content using LLMs for a robotic intervention of ASD is still an active area of research. To the best of our knowledge, the only current work that combines a social robot with using LLMs for ASD intervention is [54]. In this, the authors have used a Pepper robot integrating ChatGPT to meet the therapeutic goals of individuals with ASD. However, the structured component used for the problem-solving session is scripted, and the LLM integration is essentially an add-on to provide additional feedback and insights. Hence, it still does not fully meet the autonomy requirements for content generation during therapy sessions, which the authors pointed out in [26].
Hence, the motivation of this paper comes from the absence of a system that can be used for real-time content generation and take another step towards delivering high-end therapeutic scenarios towards a goal-oriented therapy.
III Primary Contributions
The main contributions of this paper are:
To the best of our knowledge, this is the first work that generates content to teach perspective-taking to children with ASD using LLMs.
we compare a sequence-to-sequence model (BART) with a decoder-based model (GPT-2) for text-generation tasks using BERTscore and also verify the outcome using human evaluations.
we evaluate the text generation for appropriateness for children with ASD through human evaluators from the clinical and special education domains.
IV Methodology
In this paper, the robot takes up three roles out of six possible suggested by the authors in [26] with the KASPER robot. Here, instead of the KASPER robot, we have used the NAO robot to take up the role of a provoker, reinforcer, and prompter. The workflow has been shown in Figure 2.
IV-A Role of the robot
To take up the role of the stimulator, the NAO robot generates contexts that are based on perspective-taking scenarios for the child with ASD. Then, NAO generates questions for the child to answer based on those contexts. If the child stays quiet for over 10 seconds, NAO repeats the question. If there is no reply from the child’s end, the robot gives prompts in the form of three options. This prompting also occurs when the child answers incorrectly. However, if the child can answer correctly, the robot praises the child for performing well and then moves on to a new example. This cycle continues until the termination step.
IV-B Role of the therapist
The decision of selecting the accuracy of the answers in both scenarios (one where prompts are given in the form of options and the other when no prompts are given) lies in the hands of the human instructor. The instructor is also responsible for initiating the robot, whether the answers given by the child are correct or not. Only then will the robot switch to the next pipeline segment shown in Figure 2.
A combination of the GPT and BART architectures generates these contexts, questions, and options. For this, we use the SOCIALIQA dataset, which is a social commonsense reasoning benchmark [55]. This dataset has 38,000 data points containing contexts, questions based on those contexts, and multiple-choice options for those questions. The questions broadly fall into three categories. Based on the context, the questions can inquire 1) about motivation, 2) about next steps, or 3) emotional reactions.
V Problem Formulation
As discussed, context is generated by fine-tuning GPT-2 on the SOCIALIQA dataset’s contexts. Once fined-tuned, contexts are generated using just the start and end of text prompts:
where is the generated context by fine-tuned GPT-2. For fine-tuning the BART model to generate questions and the possible answer options, BartForConditionalGeneration is used. Figure 3 shows the model that has been used in this paper. For every context vector , the BART model generates text in the following way:
where is the index of the data points, , , and, are the total number of tokens for the question, option A, option B and, option C, respectively. Further, , , and, are the predicted question, option A, option B and option C respectively. As can be seen from Figure 3, the same BART model is fine-tuned with a combination of different training inputs. Equations 2a-2d give a hierarchical nature to the training where the context is first used to predict the question and so on. For each of these predictions, the cross entropy loss is calculated and added for backpropagation:
where , , and, are the cross entropy losses calculated from predicting the question, option A, option B, and option C respectively.
This approach was also compared to the unsupervised GPT-2 generation of the question and three options based on the context. We introduce three special tokens for controlled text generation using GPT-2. In addition to the and tokens, we also use <context>:, <question>:, <ansa>:, <ansb>: and, <ansc>: to control the generation process of each of the individual parts for each data point. Once these tokens have been added to the tokenizer and the dataset tokenized, we then fine-tune the GPT-2 model and use the cross entropy loss. Each data point before tokenization can be represented as:
VI Results and Discussions
As discussed previously, we have used the SOCIALIQA dataset for all of our generation tasks. We used the premium GPU from Google Colab with A100 GPUs for training. The dataset was split into train (75%), eval (15%), and test data (10%). Adam was chosen to be the optimizer. The SOCIALIQA dataset has one context, one question, and three options for every data point. So, we fine-tuned GPT-2 for generating these contexts and then fed the output of the GPT-2 into our BART pipeline explained in the equation V to generate the questions and the options. We compared this pipeline with a similar pipeline, replacing BART with GPT-2. Further, we also present human evaluations later to validate the results and the choice of the pipeline.
VI-A Context generation using GPT-2
For generating the context, we fine-tuned a GPT-2 model based on the contexts in the SOCIALIQA dataset. To demonstrate how similar the generated contexts were to the test set, we took one thousand GPT-2 generated texts (after fine-tuning) and the same number of data points from the test set. Further, we used the T-SNE plots to visualize how the test subset and the GPT-2 generated set overlap. We vectorized the generated texts and the subset of the test dataset using scikit-learn’s TfidfVectorizer to convert both the test and the generated datasets into a tf-idf feature matrix [56]. Once vectorized, the T-SNE plot was used to visualize the high-dimensional text data into two components. Since we wanted to compare how similar the GPT-2 generated texts and the actual test set, we just used the number of components to differentiate between them. As seen from Figure 4a, the generated and the test set have a very large overlap. To characterize this overlap, we considered the data points of the scatter plot and used a Support Vector Classifier (SVC) with a radial basis kernel to see if the classifier could separate the data points into two distinct classes. Figure 4b and 4c show that the classification accuracy is 0.51, and the F1-score is around 0.55, too. This low F1-score signifies that the classifier could not very distinctly classify the values of the T-SNE plot (essentially the low-dimensional embeddings of the high-dimensional text data) [56]. The low scores on the classifying metrics signify that the text generated by fine-tuning GPT-2 generates data similar to the test set.
VI-B Evaluation metrics
We use BERTscore to evaluate our generation task [57]. We do this across generations of questions, options, and both of them combined. Other metrics are used in the literature, like the BLEU score [58]. The reason for using BERTscore is because it uses cosine similarity instead of just making an n-gram-based comparison between the actual text and the generated ones. In addition, for calculating the cosine similarities, it uses BERT embeddings between each generated token and every reference token [57]. This ensures not only exact phrase matching between the reference and the candidate text, but also the contextual relevance between the two. Since there are different BERT-based models to calculate the BERTscore from, we choose three of those models: 1) ‘distillery-base-uncased’, 2) ‘roberta-base’ and 3) ‘microsoft/deberta-large-mnlimicrosoft/deberta-xlarge-mnli’. Among these, microsoft/deberta-xlarge-mnli is currently the closest to human evaluations according to [59].
Model used for | ||||||||||
Model | BERTscore | Precision | Recall | F-1 score | Precision | Recall | F-1 score | Precision | Recall | F-1 score |
distilbert-base-uncased |
0.78 | 0.79 | 0.79 | 0.75 | 0.75 | 0.75 | 0.84 | 0.85 | 0.85 | |
BART |
roberta-base |
0.88 | 0.88 | 0.88 | 0.76 | 0.77 | 0.77 | 0.86 | 0.86 | 0.86 |
mnlimicrosoft/deberta-xlarge-mnli |
0.62 | 0.61 | 0.61 | 0.57 | 0.54 | 0.55 | 0.68 | 0.66 | 0.67 | |
distilbert-base-uncased |
0.71 | 0.76 | 0.73 | 0.70 | 0.74 | 0.72 | 0.80 | 0.83 | 0.81 | |
GPT-2 |
roberta-base |
0.85 | 0.87 | 0.86 | 0.84 | 0.85 | 0.84 | 0.90 | 0.91 | 0.90 |
mnlimicrosoft/deberta-xlarge-mnli |
0.53 | 0.57 | 0.55 | 0.49 | 0.53 | 0.51 | 0.69 | 0.71 | 0.70 |
VI-C Ablation study
This section compares using a BART pipeline for question and options generation vs an analogous GPT -2-based generation task. We evaluate the BERTscore for generated questions, options, and both for each test sample. As seen from Table I, our BART pipeline performs better on average for all of the generation tasks, at least when we use the Microsoft/deberta-xlarge-mnli model for calculating the BERTscore. This result contributes to our choice of using the BART pipeline because we use a Masked Language Modeling approach for it. Using masks makes this pipeline a supervised learning approach, which gives control over text generation.
Context |
Reference |
BART |
GPT-2 |
After several days of arguing with the boss, Bailey finally was exhausted. Bailey let it be. | Question: | Question: | Question: |
How would Bailey feel afterwards? |
What will Bailey want to do next? |
What will Bailey want to do next? |
Options: | Options: | Options: | |
A: ignored and defeated |
A: go home and rest |
A: sleep |
B: amenable |
B: get a new job |
B: break a sweat |
C: happy that the boss reconsidered |
C: go back to work |
C: quit the battle |
Cohen’s kappa (weighted) for BART pipeline |
R1 and R2 (0.705) |
R1 and R2 (0.639) |
R1 and R4 (0.693) |
R1 and R4 (0.787) |
R1 and R3 (0.72) |
R2 and R6 (0.680) |
R1 and R5 (0.85) |
R1 and R4 (0.70) |
R1 and R6 (0.80) |
R1 and R6 (0.634) |
R2 and R6 (0.85) |
R2 and R3 (0.715) |
R4 and R5 (0.87) |
R3 and R6 (0.8344) |
Cohen’s kappa (weighted) for GPT-2 pipeline | |||
R2 and R6 (0.64) |
An example comparing the BART and the GPT-2 pipeline for question and options generation can be seen in Table II. For a context taken from the SOCIALIQA dataset, we generate the text, evaluate it with the reference question and options from the dataset and calculate the score. We take five random examples from the SOCIALIQA dataset and ask six raters to score them between 0-10. Cohen’s kappa (weighted) was calculated to check for inter-rater reliability for each generation task (separately for question generation, options generation, and question + options generation) for each of the BART and GPT-3 pipelines. Table III shows the kappa values between the raters. We have only considered those values between the raters with a kappa higher than 0.6 (moderate agreement). Based on this threshold, the kappa values for the GPT-2 pipeline were only higher than 0.6 for rating . This low value can be attributed to fewer examples (five) considered for ratings. For comparing the BART and GPT pipelines based on the human evaluations for the whole text generation pipeline (), Figure 5 compares the values of the BERTscore with human evaluations for BART and GPT-2.
As evident from Figure 5, the BART pipeline performs better on the BERTscore metric, which is also consistent with human-evaluations. The human evaluation scores are very low for GPT-2 compared to the BERTscores since the human evaluations were performed on a very small sample of data with six raters.
VI-D Appropriateness for children with ASD
We finally concatenated our fine-tuned GPT-2 model from Section VI-A with the BART pipeline to generate the context, question, and options. An example from the generated texts for our GPT-2 + BART pipeline has been shown in Table IV. These texts were rated by human evaluators on two parameters:
Context |
Question |
Options |
Addison made Sydney’s bed for them. |
What will Sydney want to do next? |
A. thank Addison for helping |
B. get a new bed |
C. go to the store |
Q1: Is the social situation appropriate for children with ASD between the age range of 8-12 years (0-10 scale)?
Q2: Is there at least one appropriate answer from the options generated by the robot (0 or 1 scale)?
The scores for the first question were divided into binary scale (rater scores 0-7: 0 and 8-10: 1) for the inter-rater agreement between raters and 0-10 scale for finding the average of the scores. Four raters (R1, R2, R3, and R6) from Table III have clinical or teaching experience with children with ASD. The highest Cohen’s kappa was found between R2 and R6 (0.68). So, we calculated the average scores of these raters and the average score for Q1 was found to be 6.8 on a scale of 0-10. In addition, we found the percentage agreement between the raters for Q2 to be 0.95.
VII Conclusion
In this paper, we attempt a scalable solution for a robotic intervention of children with ASD. We use an LLM pipeline with GPT-2 and BART to generate texts that the robot vocalises and can be used in a clinical setting in the presence of experts. The robot in this paper has been given the roles of a stimulator, prompter, and reinforcer, out of which the stimulator roles have autonomy in the text generation the robot uses. The automatic text generation is evaluated not only through the BERTscore metric but also through human evaluators. In addition, the appropriateness of the generated social situation using our LLM pipeline is also scored by experts with backgrounds in clinical intervention or teaching children with ASD.
As mentioned in the earlier sections, this work is the first of its kind to develop such a robot to automatically generate content during an intervention scenario for children with ASD. So, the scope of future work is immense. This system can be made more versatile in scenarios to be taught. For example, the LLM pipeline can be fine-tuned to generate content specific to a school setting, home setting, or even for making natural conversations, depending on the dataset used for fine-tuning. Therefore, the direction of our future work too lies in expanding the number of social situations for children with ASD and adapting to their specific needs during the intervention.